# Polylog-time- and constant-space-overhead fault-tolerant quantum computation with quantum low-density parity-check codes

### v1

@@ VI.2 Partial circuit reduction
$\leq\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{|A^{\prime}|}\sum_{k=0}^{r}\binom{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}{k}\binom{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}{r-k}\quantity(p^{(1)})^{k}\quantity(p^{(2)})^{r-k}$ (254)
$\leq\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{|A^{\prime}|}\quantity(p^{(2)})^{r}\sum_{k=0}^{r}\binom{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}{k}\binom{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}{r-k}\quantity(\frac{p^{(1)}}{p^{(2)}})^{k}.$ (255)
We use the inequality $\binom{n}{r-k}\leq 2^{n}$ for $\binom{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}{r-k}$ to obtain
$\binom{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}{r-k}$ $\leq 2^{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}.$ (256)
Then, we extend the summation over $k$ to the range from 00 to $c_{\mathrm{bw}}^{(1)}|A^{\prime}|$, we have
$\sum_{k=0}^{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}\binom{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}{k}\left(\frac{p^{(1)}}{\beta}\right)^{k}=\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}.$ (257)
Thus, we obtain
$\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{|A^{\prime}|}\sum_{k=0}^{r}\binom{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}{k}\binom{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}{r-k}\quantity(p^{(1)})^{k}\quantity(p^{(2)})^{r-k}$ (258)
$\leq 2^{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{|A^{\prime}|}\quantity(p^{(2)})^{r}.$ (259)
Since the case of $p^{(2)}=1$ is trivial, we consider the case of $p^{(2)}<1$ so the sum over $r$ is bounded by
$\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{|A^{\prime}|}\quantity(p^{(2)})^{r}$ $\leq\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{\infty}\quantity(p^{(2)})^{r}=\frac{\quantity(p^{(2)})^{\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}}{1-p^{(2)}}$ (260)
$\leq\frac{\quantity(p^{(2)})^{\frac{|A^{\prime}|}{c_{\mathrm{fw}}}}}{1-p^{(2)}}.$ (261)
Finally, we have
$\mathbb{P}[\Gamma(F)\supseteq A^{\prime}]$ $\leq 2^{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}\frac{\quantity(p^{(2)})^{\frac{|A^{\prime}|}{c_{\mathrm{fw}}}}}{1-p^{(2)}},$ (262)
which is upper bounded by
$\mathbb{P}[\Gamma(F)\supseteq A^{\prime}]$ $\leq\left(2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}}\frac{\quantity(p^{(2)})^{1/c_{\mathrm{fw}}}}{1-p^{(2)}}\right)^{|A^{\prime}|}.$ (263)
By condition 4, we have
$\quad\forall j\in C^{\prime},\quad q_{j}\geq 2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}}\frac{\quantity(p^{(2)})^{1/c_{\mathrm{fw}}}}{1-p^{(2)}}.$ (264)
Therefore,
$\mathbb{P}[\Gamma(F)\supseteq A^{\prime}]\leq\prod_{j\in A^{\prime}}q_{j}.$ (265)
@@ VI.3 Proof of existence of threshold
$\omega 2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)q_{\mathrm{prep}}^{\mathrm{th}})^{1/2^{d_{0}}}=p^{\mathrm{dec}}_{\mathrm{th}}.$ (266)
@@
$\omega 2^{2d_{0}}\quantity(1+\frac{dq_{\mathrm{gate}}^{\mathrm{th}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}=p_{\mathrm{dec}}^{\mathrm{th}},$ (267)
$\omega 2^{2^{d_{0}}}\quantity(1+\frac{dq_{\mathrm{meas}}^{\mathrm{th}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}=p_{\mathrm{dec}}^{\mathrm{th}},$ (268)
@@
Lemma 17 (Circuit reducibility of rectagles).
@@
is $\delta_{3}(n,\varepsilon)$-reducible to $\quantity(C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}},\{p^{\prime}_{j}\}_{j\in C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}}})$, where ${C^{\prime\prime}}^{\mathrm{wait}}$ is a single-depth circuit of wait locations, and $\{p^{\prime}_{j}\}$ is the set of error parameters given by
@@
Here, $M\geq 1$ is a constant. Since ${C^{\prime\prime}}^{\mathrm{wait}}\cup C_{\mathrm{sync}}^{\mathrm{wait}}$ consists only of wait operations, it is obviously reducible to ${C^{\prime\prime}}^{\mathrm{wait}}$ with error parameter of each wait location bounded by $(2d+M)p_{\mathrm{loc}}$. Thus, from Lemmas 14 and 15, $\{C,\{p_{i}\}_{i\in C}\}$ is $\delta_{3}(n,\varepsilon)$-reducible to an incomplete circuit
$C^{\prime\prime}\coloneqq C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}}\cup C^{\mathrm{EC}},$ (300)
@@
$p^{\prime}_{j}$ $=0$ if $\text{if $j\in C^{\mathrm{prep}}$},$ (301)
$p^{\prime}_{j}$ $\leq(2d+M)p_{\mathrm{loc}}$ if $\text{if $j\in{C^{\prime\prime}}^{\mathrm{wait}}$},$ (302)
$p^{\prime}_{j}$ $\leq p_{\mathrm{loc}}$ if $\text{if $j\in{C}^{\mathrm{EC}}$}.$ (303)
@@
${C^{\prime}}^{\mathrm{synd}}\coloneqq{C^{\prime\prime}}^{\mathrm{wait}}\cup C^{\mathrm{synd}},$ (304)
@@
$p^{\prime}_{i}$ $\leq(2d+M)p_{\mathrm{loc}}$ if $\text{if $i\in{C^{\prime}}^{\mathrm{synd}}$}.$ (305)
@@
For the reduction, we consider error propagation from a given set of faulty locations in ${C^{\prime}}^{\mathrm{synd}}$ to the set of faulty locations by propagating Pauli errors until the time step after the $Z$-basis measurements or the final time step of the wait locations in ${C^{\prime}}^{\mathrm{synd}}$. Let us denote the non-faulty locations in the middle of ${C}^{\prime\mathrm{synd}}$ after this error propagation by ${C}^{\prime\mathrm{synd}}_{\mathrm{nonfaulty}}$, and the faulty locations by ${C}^{\prime\mathrm{synd}}_{\mathrm{nonfaulty}}$, i.e.,
@@
$|\Gamma^{\mathrm{synd}}(\{i\})|\leq 2^{d_{0}}\text{~{}for all~{}}i\in{C^{\prime}}^{\mathrm{synd}}.$ (307)
Moreover, in the error propagation, each location in ${C}_{\mathrm{faulty}}^{\mathrm{synd}}$ may be affected by errors occurring in at most $(d_{0}+1)\times 2^{d_{0}}$ possible locations in ${C^{\prime}}^{\mathrm{synd}}$; thus we have
|$|\{i\in{C^{\prime}}^{\mathrm{synd}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|\leq(d_{0}+1)2^{d_{0}}$ (308)
 for all $\text{~{}for all~{}}j\in{C^{\prime}}^{\mathrm{synd}}.$ (309)
As a result, if we choose
$p^{\prime\prime}_{j}=\omega 2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)p_{\mathrm{loc}})^{1/2^{d_{0}}},$ (310)
for all $\text{for all $j\in{C}^{\prime\mathrm{synd}}_{\mathrm{faulty}}$},$ (311)
with $\omega$ as defined in (162), the map $\Gamma^{\mathrm{synd}}$ and the sets of error parameters $\{p^{\prime}_{i}\}_{i\in{C^{\prime}}^{\mathrm{synd}}}$ and $\{p^{\prime\prime}_{j}\}_{j\in{C}^{\prime\mathrm{synd}}_{\mathrm{faulty}}}$ satisfy all conditions in Lemma 16. Thus, ${C^{\prime}}^{\mathrm{synd}}$ is reducible from $\{p^{\prime}_{i}\}_{i}$ to $\{p^{\prime\prime}_{j}\}_{j}$.
@@
$p_{\mathrm{phys}}$ $\leq\omega 2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)p_{\mathrm{loc}})^{1/2^{d_{0}}},$ (312)
@@
$p_{\mathrm{loc}}<\frac{1}{2d+M}\quantity(\frac{p_{\mathrm{dec}}^{\mathrm{th}}}{\omega 2^{(d_{0}+1)2^{d_{0}}}})^{2^{d_{0}}}=q^{\mathrm{th}}_{\mathrm{prep}}$ (314)
@@
$\omega 2^{(d_{0}+1)2^{d_{0}}}\quantity(\omega(2d+M)p_{\mathrm{loc}})^{1/2^{d_{0}}}<p^{\mathrm{dec}}_{\mathrm{th}},$ (315)
@@
For the reduction, we consider error propagation from a given set of faulty locations in ${C^{\prime\prime}}$ to the set of locations until the time step after the $Z$-basis measurements or the final time step of the wait locations in ${C^{\mathrm{synd}}}$. Let us denote the non-faulty locations in the middle of $C^{\prime\prime}$ after this error propagation by $C^{\prime\prime}_{\mathrm{nonfaulty}}$, and the faulty locations by $C^{\prime\prime}_{\mathrm{nonfaulty}}$, i.e.,
@@
$|\Gamma(\{i\})|\leq 2^{2d_{0}}\text{~{}for all~{}}i\in{C^{\prime}}^{\mathrm{synd}}.$ (331)
@@
|$|\{i\in{C}^{\mathrm{wait}}_{\mathrm{re}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|\leq 2^{2d_{0}},$ (332)
|$|\{i\in{C^{\prime\prime}}\setminus{C}^{\mathrm{wait}}_{\mathrm{re}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|\leq 2^{2d_{0}+1}(d_{0}+1).$ (333)
As a result, if we choose
$p^{\prime\prime}_{j}=\omega 2^{2^{2d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}},$ (334)
for all $\text{for all $j\in C^{\prime\prime}_{\mathrm{faulty}}$},$ (335)
with $\omega$ as defined in (162), the map $\Gamma$ and the sets of error parameters $\{p_{i}\}_{i\in{C^{\prime\prime}}}$ and $\{p^{\prime\prime}_{j}\}_{j\in C^{\prime\prime}_{\mathrm{faulty}}}$ satisfy all conditions in Lemma 16. Thus, a pair $\quantity(C^{\prime\prime},\{p_{i}\}_{i\in C^{\prime\prime}})$ is reducible to $\{C^{\prime\prime}_{\mathrm{faulty}},\{p^{\prime\prime}_{j}\}_{j\in C^{\prime\prime}_{\mathrm{faulty}}}\}$.
@@
$p_{\mathrm{phys}}$ $\leq\omega 2^{2d_{0}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}},$ (336)
$p_{\mathrm{synd}}$ $\leq\omega 2^{2d_{0}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}.$ (337)
@@
$\omega 2^{2d_{0}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}<p_{\mathrm{dec}}^{\mathrm{th}},$ (339)
@@
Let us consider circuit reduction of an incomplete circuit $C$. For the reduction, we consider the propagation of Pauli errors at the faulty locations in $C$ to the time step after the $Z$-basis measurements in ${C}^{\mathrm{meas}}$. Let us denote the non-faulty locations in the middle of $C$ after this error propagation by $C_{\mathrm{nonfaulty}}$, and the faulty locations by $C_{\mathrm{faulty}}$, i.e.,
@@
$|\Gamma(\{i\})|\leq 2^{d_{0}}\text{~{}for all~{}}i\in{C^{\prime}}^{\mathrm{synd}}.$ (349)
@@
|$|\{i\in{C}^{\mathrm{wait}}_{\mathrm{re}}\colon j\in\Gamma(\{i\})\}|\leq 2^{d_{0}},$ (350)
|$|\{i\in C^{\mathrm{wait}}\cup C^{\mathrm{meas}}\colon j\in\Gamma(\{i\})\}|\leq 2^{d_{0}+1}\times(d_{0}+1).$ (351)

### v2

@@ VI.2 Partial circuit reduction
$\leq\mathbb{P}[|F\cap\Lambda_{1}(A^{\prime})|\geq k_{0}]+$
$~{}~{}~{}~{}\mathbb{P}[|F\cap\Lambda(A^{\prime})|\geq k_{0}\text{~{}and~{}}|F\cap\Lambda_{1}(A^{\prime})|\leq k_{0}]$ (273)
$\leq\left(2^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(1)})^{1/c_{\mathrm{fw}}}+2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(2)})^{1/c_{\mathrm{fw}}}\right)^{|A^{\prime}|}$ (274)
$\leq\prod_{j\in A^{\prime}}q_{j}$ (275)
from condition 4 with $A^{\prime}\subseteq C_{\mathrm{faulty}}^{\prime}$ and $|A^{\prime}|\geq 1$, which proves (247).
@@ VI.3 Proof of existence of threshold
$2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)q_{\mathrm{prep}}^{\mathrm{th}})^{1/2^{d_{0}}}=p^{\mathrm{dec}}_{\mathrm{th}}.$ (276)
@@
$2^{2^{2d_{0}}}\quantity(1+\frac{dq_{\mathrm{gate}}^{\mathrm{th}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{2d_{0}+1}(d_{0}+1)}(dq_{\mathrm{gate}}^{\mathrm{th}})^{1/2^{2d_{0}}}=p_{\mathrm{dec}}^{\mathrm{th}},$ (277)
$2^{2^{d_{0}}}\quantity(1+\frac{dq_{\mathrm{meas}}^{\mathrm{th}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{d_{0}+1}(d_{0}+1)}(dq_{\mathrm{meas}}^{\mathrm{th}})^{1/2^{d_{0}}}=p_{\mathrm{dec}}^{\mathrm{th}}$ (278)
@@
Lemma 17 (Circuit reducibility of rectangles).
@@
is $\delta_{3}$-reducible to $\quantity(C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}},\{p^{\prime}_{j}\}_{j\in C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}}})$, where ${C^{\prime\prime}}^{\mathrm{wait}}$ is a single-depth circuit of wait locations, and $\{p^{\prime}_{j}\}$ is the set of error parameters given by
@@
Here, $M\geq 1$ is a constant. Since ${C^{\prime\prime}}^{\mathrm{wait}}\cup C_{\mathrm{sync}}^{\mathrm{wait}}$ consists only of wait operations, it is obviously reducible to ${C^{\prime\prime}}^{\mathrm{wait}}\cup C_{\mathrm{sync}}^{\mathrm{wait}}$ with the error parameter at each wait location in $C_{\mathrm{sync}}^{\mathrm{wait}}$ bounded by $(2d+M)p_{\mathrm{loc}}$, while the error parameter of each wait location in ${C^{\prime\prime}}^{\mathrm{wait}}$ is zero. Thus, from Lemmas 14 and 15, $\{C,\{p_{i}\}_{i\in C}\}$ is $\delta_{3}$-reducible to an incomplete circuit
$C^{\prime\prime}\coloneqq C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}}\cup{C}^{\mathrm{wait}}_{\mathrm{sync}}\cup C^{\mathrm{EC}},$ (310)
@@
$p^{\prime}_{j}$ $=0$ if $\text{if $j\in C^{\mathrm{prep}}\cup{C^{\prime\prime}}^{\mathrm{wait}}$},$ (311)
$p^{\prime}_{j}$ $\leq(2d+M)p_{\mathrm{loc}}$ if $\text{if $j\in{C}^{\mathrm{wait}}_{\mathrm{sync}}$},$ (312)
$p^{\prime}_{j}$ $\leq p_{\mathrm{loc}}$ if $\text{if $j\in{C}^{\mathrm{EC}}$},$ (313)
where
@@
${C^{\prime}}^{\mathrm{synd}}\coloneqq{C}^{\mathrm{wait}}_{\mathrm{sync}}\cup{C^{\prime\prime}}^{\mathrm{wait}}\cup C^{\mathrm{synd}},$ (314)
@@
$p^{\prime}_{j}$ $=0$ if $i\in{C}^{\mathrm{wait}}_{\mathrm{sync}}$ (315)
$p^{\prime}_{i}$ $\leq(2d+M)p_{\mathrm{loc}}$ if $\text{if $i\in{C^{\prime\prime}}^{\mathrm{wait}}\cup C^{\mathrm{synd}}$}.$ (316)
@@
To apply Lemma 16 for the circuit reduction of $C^{\prime\mathrm{synd}}$, we consider a partition of the set $C^{\prime\mathrm{synd}}$ into disjoint sets $C_{1}$ and $C_{2}$. Here, $C_{1}$ corresponds to ${C}^{\mathrm{wait}}_{\mathrm{sync}}$, and $C_{2}$ corresponds to ${C^{\prime\prime}}^{\mathrm{wait}}\cup C^{\mathrm{synd}}$. We consider error propagation from a given set of faulty locations in ${C^{\prime}}^{\mathrm{synd}}$ to the set of faulty locations by propagating Pauli errors until the time step after the $Z$-basis measurements or the final time step of the wait locations in ${C^{\prime}}^{\mathrm{synd}}$. Let us denote the non-faulty locations in the middle of ${C}^{\prime\mathrm{synd}}$ after this error propagation by ${C}^{\prime\mathrm{synd}}_{\mathrm{nonfaulty}}$, and the faulty locations by ${C}^{\prime\mathrm{synd}}_{\mathrm{nonfaulty}}$, i.e.,
@@
$|\Gamma^{\mathrm{synd}}(\{i\})|\leq c_{\mathrm{fw}}=2^{d_{0}}\text{~{}for all~{}}i\in{C^{\prime}}^{\mathrm{synd}}.$ (318)
Moreover, in the error propagation, each location in ${C}_{\mathrm{faulty}}^{\mathrm{synd}}$ may be affected by errors occurring in at most $(d_{0}+1)\times 2^{d_{0}}$ possible locations in ${C^{\prime}}^{\mathrm{synd}}$. Thus we have for all $j\in{C}^{\prime\mathrm{synd}}_{\mathrm{faulty}}$
|$|\{i\in C_{\mathrm{sync}}^{\mathrm{wait}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|\leq c_{\mathrm{bw}}^{(1)}=2^{d_{0}},$ (319)
|$|\{i\in{C^{\prime\prime}}^{\mathrm{wait}}\cup C^{\mathrm{synd}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|$
$\leq c_{\mathrm{bw}}^{(2)}=(d_{0}+1)2^{d_{0}}.$ (320)
Furthermore, under the partition, we have $p_{i}=p^{(1)}=0$ for all $i\in C^{\mathrm{wait}}_{\mathrm{sync}}$ and $p_{i}\leq p^{(2)}=(2d+M)p_{\mathrm{loc}}$ for all $i\in C^{\prime\prime\mathrm{wait}}\cup C^{\mathrm{synd}}$. As a result, if we choose $p^{\prime\prime}_{j}$ for all $j\in{C}^{\prime\mathrm{synd}}_{\mathrm{faulty}}$ as
$p^{\prime\prime}_{j}$ $=2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(2)})^{1/c_{\mathrm{fw}}}+2^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(1)})^{1/c_{\mathrm{fw}}}$ (321)
$=2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)p_{\mathrm{loc}})^{1/2^{d_{0}}},$ (322)
the map $\Gamma^{\mathrm{synd}}$ and the sets of error parameters $\{p^{\prime}_{i}\}_{i\in{C^{\prime}}^{\mathrm{synd}}}$ and $\{p^{\prime\prime}_{j}\}_{j\in{C}^{\prime\mathrm{synd}}_{\mathrm{faulty}}}$ satisfy all conditions in Lemma 16. Thus, $\quantity({C^{\prime}}^{\mathrm{synd}},\{p^{\prime}_{i}\}_{i\in{C^{\prime}}^{\mathrm{synd}}})$ is reducible to $\quantity({C^{\prime}}^{\mathrm{synd}}_{\mathrm{faulty}},\{p^{\prime\prime}_{j}\}_{j\in{C^{\prime}}^{\mathrm{synd}}_{\mathrm{faulty}}})$.
@@
$p_{\mathrm{phys}}$ $\leq 2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)p_{\mathrm{loc}})^{1/2^{d_{0}}},$ (323)
@@
$p_{\mathrm{loc}}<\frac{1}{2d+M}\quantity(\frac{p_{\mathrm{dec}}^{\mathrm{th}}}{2^{(d_{0}+1)2^{d_{0}}}})^{2^{d_{0}}}=q^{\mathrm{th}}_{\mathrm{prep}}$ (325)
@@
$2^{(d_{0}+1)2^{d_{0}}}\quantity((2d+M)p_{\mathrm{loc}})^{1/2^{d_{0}}}<p^{\mathrm{dec}}_{\mathrm{th}},$ (326)
@@
To apply Lemma 16 for the reduction of $C^{\prime\prime}$, we consider a partition of the set $C^{\prime\prime}$ into disjoint sets $C_{1}$ and $C_{2}$. Here, $C_{1}$ corresponds to $C^{\prime\prime}\setminus C_{\mathrm{re}}^{\mathrm{wait}}$, and $C_{2}$ corresponds to $C_{\mathrm{re}}^{\mathrm{wait}}$. We consider error propagation from a given set of faulty locations in ${C^{\prime\prime}}$ to the set of locations until the time step after the $Z$-basis measurements or the final time step of the wait locations in ${C^{\mathrm{synd}}}$. Let us denote the non-faulty locations in the middle of $C^{\prime\prime}$ after this error propagation by $C^{\prime\prime}_{\mathrm{nonfaulty}}$, and the faulty locations by $C^{\prime\prime}_{\mathrm{nonfaulty}}$, i.e.,
@@
$|\Gamma(\{i\})|\leq c^{\prime}_{\mathrm{fw}}=2^{2d_{0}}\text{~{}for all~{}}i\in{C^{\prime}}^{\mathrm{synd}}.$ (342)
@@
|$|\{i\in{C^{\prime\prime}}\setminus{C}^{\mathrm{wait}}_{\mathrm{re}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|$
$\leq c_{\mathrm{bw}}^{\prime(1)}=2^{2d_{0}+1}(d_{0}+1),$ (343)
|$|\{i\in{C}^{\mathrm{wait}}_{\mathrm{re}}\colon j\in\Gamma^{\mathrm{synd}}(\{i\})\}|\leq c_{\mathrm{bw}}^{\prime(2)}=2^{2d_{0}}.$ (344)
Furthermore, under the partition, we have $p_{i}=p^{(1)}=dp_{\mathrm{loc}}$ for all $i\in C^{\prime\prime}\setminus C^{\mathrm{wait}}_{\mathrm{re}}$ and $p_{i}\leq p^{(2)}=\lambda$ for all $i\in C^{\mathrm{wait}}_{\mathrm{re}}$. As a result, if we choose $p^{\prime\prime}_{j}$ for all $j\in C^{\prime\prime}_{\mathrm{faulty}}$ as
$p^{\prime\prime}_{j}$ $=2^{c_{\mathrm{bw}}^{\prime(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{\prime(1)}}\quantity(p^{(2)})^{1/c^{\prime}_{\mathrm{fw}}}+2^{c_{\mathrm{bw}}^{\prime(1)}}\quantity(p^{(1)})^{1/c^{\prime}_{\mathrm{fw}}}$ (345)
$=2^{2^{2d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{2d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{2d_{0}}},$ (346)
the map $\Gamma$ and the sets of error parameters $\{p_{i}\}_{i\in{C^{\prime\prime}}}$ and $\{p^{\prime\prime}_{j}\}_{j\in C^{\prime\prime}_{\mathrm{faulty}}}$ satisfy all conditions in Lemma 16. Thus, a pair $\quantity(C^{\prime\prime},\{p_{i}\}_{i\in C^{\prime\prime}})$ is reducible to $\{C^{\prime\prime}_{\mathrm{faulty}},\{p^{\prime\prime}_{j}\}_{j\in C^{\prime\prime}_{\mathrm{faulty}}}\}$.
@@
$p_{\mathrm{phys}}$ $\leq 2^{2^{2d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{2d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{2d_{0}}},$ (347)
$p_{\mathrm{synd}}$ $\leq 2^{2^{2d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{2d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{2d_{0}}}.$ (348)
@@
$2^{2^{2d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{2d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{2d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{2d_{0}}}<p_{\mathrm{dec}}^{\mathrm{th}},$ (350)
@@
To apply Lemma 16 for the reduction of $C$, we consider a partition of the set $C$ into disjoint sets $C_{1}$ and $C_{2}$. Here, $C_{1}$ corresponds to $C\setminus C_{\mathrm{re}}^{\mathrm{wait}}$, and $C_{2}$ corresponds to $C_{\mathrm{re}}^{\mathrm{wait}}$. We consider the propagation of Pauli errors at the faulty locations in $C$ to the time step after the $Z$-basis measurements in ${C}^{\mathrm{meas}}$. Let us denote the non-faulty locations in the middle of $C$ after this error propagation by $C_{\mathrm{nonfaulty}}$, and the faulty locations by $C_{\mathrm{faulty}}$, i.e.,
@@
$|\Gamma(\{i\})|\leq c^{\prime\prime}_{\mathrm{fw}}=2^{d_{0}}\text{~{}for all~{}}i\in{C^{\prime}}^{\mathrm{synd}}.$ (360)
@@
|$|\{i\in C^{\mathrm{wait}}\cup C^{\mathrm{meas}}\colon j\in\Gamma(\{i\})\}|\leq c_{\mathrm{bw}}^{\prime\prime(1)}=2^{d_{0}+1}(d_{0}+1),$ (361)
|$|\{i\in{C}^{\mathrm{wait}}_{\mathrm{re}}\colon j\in\Gamma(\{i\})\}|\leq c_{\mathrm{bw}}^{\prime\prime(2)}=2^{d_{0}}.$ (362)
Furthermore, under the partition, we have $p_{i}=p^{(1)}=dp_{\mathrm{loc}}$ for all $i\in C\setminus C^{\mathrm{wait}}_{\mathrm{re}}$ and $p_{i}\leq p^{(2)}=\lambda$ for all $i\in C^{\mathrm{wait}}_{\mathrm{re}}$. As a result, if we choose $p^{\prime}_{j}$ for all $j\in C_{\mathrm{faulty}}$ as
$p^{\prime}_{j}$ $=2^{c_{\mathrm{bw}}^{\prime\prime(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{\prime\prime(1)}}\quantity(p^{(2)})^{1/c^{\prime\prime}_{\mathrm{fw}}}+2^{c_{\mathrm{bw}}^{\prime\prime(1)}}\quantity(p^{(1)})^{1/c^{\prime\prime}_{\mathrm{fw}}}$ (363)
$=2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{d_{0}}},$ (364)